The 1-year risk free rate is 5.0% and there are no storage costs. If the spot price of gold is $390 and the 1-year forward price is $409.50, your arbitrage profits by selling gold futures is closest to?ause:
Answer #1

Future value = spot*e^(r*t)

=390*e^(0.05*1)

=410

Profit = futures price-forward price = 410-409.5=0.5